# Accelerating Eigenvalue Dataset Generation via Chebyshev Subspace Filter

### v1

@@ 1 Introduction
Despite their success, data-driven approaches face a fundamental limitation: the reliance on labeled datasets. Training neural networks requires large-scale labeled data, which is often generated using computationally expensive traditional methods. It usually takes dozens of hours or even days. For example, the QM9 dataset (Ramakrishnan et al., 2014) contains $1.34\times 10^{5}$ molecular data points, each produced by solving Hamiltonian operator eigenvalue problems. These calculations typically employ traditional algorithms, whose computational costs can escalate dramatically with increasing problem complexity, like finer grid resolutions or higher accuracy requirements. This scalability issue represents a significant bottleneck for generating the labeled data needed to train deep learning models. Furthermore, the diversity of scientific problems leads to the need for a unique dataset for each scenario, which further intensifies this challenge of computational intractability. As a result, the high computational expense of generating eigenvalue data severely limits the application of data-driven approaches (Zhang et al., 2023).
@@
  • To the best of our knowledge, SCSF is the first method to accelerate the operator eigenvalue data generation.
@@ 2.1 Discretization of Eigenvalue Problem
Our main focus is on solving the matrix eigenvalue problem, the most time-consuming part of eigenvalue data generation. As shown in Figure 1 (left), these problems are typically solved by numerical discretization methods such as FDM (Strikwerda, 2004; LeVeque, 2002). These discretization techniques embed the infinite-dimensional Hilbert space of operators into an appropriate finite-dimensional space, thereby transforming operator eigenvalue problems into matrix eigenvalue problems. We provide a simple example to clarify the discussed processes. A detailed process can be found in Appendix C. Specifically, we discuss the case that uses FDM to solve the eigenvalue problem of the two-dimensional Poisson operator, transforming it into a matrix eigenvalue problem:
@@
Typically, training a neural network requires a number of data from $10^{3}$ to $10^{5}$ (Lu et al., 2019). Such a multitude of eigenvalue systems, derived from the same distribution of operators, naturally exhibit a highly similarity (Soodhalter et al., 2020). It is precisely this similarity that is key to the effective acceleration of SCSF. We can conceptualize this as the task of solving a sequential series of matrix eigenvalue problems:
$A^{(i)}v^{(i)}_{j}=\lambda^{(i)}_{j}v^{(i)}_{j},\quad j=1,\cdots,L;\quad i=1,2,\cdots,N,$ (3)
@@ 4 $\sigma_{i+1}=1/(2/\sigma_{1}-\sigma_{i})$;
Chebyshev filtered subspace iteration is closely related to Chebyshev orthogonal polynomials (Mason & Handscomb, 2002; Rivlin, 2020). Chebyshev polynomials are widely used due to their strong approximation capabilities. The Chebyshev polynomials $C_{m}(t)$ of degree $m$ are defined on the interval [1,1][-1,1] and are expressed as
@@ 3.1 The Sorting Algorithm
To benefit the successive solving sequence of the eigenvalue problem, we need a sorting algorithm that pulls matrices with similar spectral properties, like invariant subspaces, close enough in the solving sequence, so that solving the current matrix in sequence can be easily boosted by the previous solving. Recalling Section 2.1, eigenvalue problem, the matrix $A^{(i)}$, is generated from the parameter matrix $P^{(i)}$ (Lu et al., 2022; Li et al., 2020). A naive strategy is to use the Frobenius distance of the parameter matrices $P^{(i)}$ to perform a greedy sort (Wang et al., 2024). And by repeatedly fetching without reservation from the remaining matrix in the dataset, we can reorganize the solving sequence so that the successive solving can benefit from the re-ordered sequence.
@@ 12  Remove $j_{min}$ from $seq_{0}$, append $j_{min}$ to ${seq}_{mat}$ and set $i_{0}=j_{min}$;
As shown in Algorithm 2, suppose we have $N$ eigenvalue problems, the parameter matrices $P^{(i)}\in\mathds{C}^{p\times p}$, and the low-frequency truncated matrices $P^{(i)}_{low}\in\mathds{C}^{p_{0}\times p_{0}}$. The computational complexity of directly using a greedy algorithm is $\mathcal{O}(N^{2}p^{2})$. Our sorting algorithm’s complexity consists of two main parts: 1. FFT Computation: The complexity of FFT is $\mathcal{O}(p^{2}\log p)$ per matrix. For $N$ matrices, this totals $\mathcal{O}(Np^{2}\log p)$. 2. Greedy Sorting: The subsequent greedy sorting algorithm has a complexity of $\mathcal{O}(N^{2}p_{0}^{2})$. Overall, the total complexity is $\mathcal{O}(N^{2}p_{0}^{2}+Np^{2}\log p)$. Since $p_{0}\ll p$ and $p\ll N$, our sorting algorithm effectively reduces computational cost.
@@ 3.2 Chebyshev Filtered Subspace Iteration
Algorithm 3 outlines the process of ChFSI for solving the $i$-th eigenvalue problem $A^{(i)}$ ($1<i\leqslant N$) where $L$ eigenvalues need to be solved. The initial approximate invariant subspace $V^{(i-1)}$ and spectral distribution $\Lambda^{(i-1)}$ are derived from the eigenvectors and eigenvalues of the previous problem $A^{(i-1)}$ in the sequence. The parameter $m$ denotes the polynomial degree in the filter function, e.g., $m=20$. For the first eigenvalue problem $A^{(1)}$ in the sequence, the initial iterative subspace $\tilde{V}_{0}$ and initial spectrum $\tilde{\Lambda}_{0}$ are randomly generated.
@@ 4.2 Main Experiment
Table 1: Comparison of average computation times (in seconds) for eigenvalue problems using various algorithms. The first row lists different algorithms, the first column details the datasets, including matrix dimensions and solution precisions (relative residual), and the second column shows the number of eigenvalues $L$ computed for each matrix. The best algorithm is in bold. The symbol ’-’ denotes the result of a method that fails to converge under the given setting.
@@ 4.3 Efficacy of Chebyshev Subspace Filter
Table 2: Impact of initial subspace modifications on average computation time (in seconds) for different algorithms. ’*’ denotes the modified version. The first row lists algorithms, and the first column shows the number of eigenvalues $L$ computed. The best algorithm is in bold, and ’-’ indicates the result of a method that fails to converge under the given setting.
@@
To analyze the efficacy of the Chebyshev subspace filter, we conducted the following experiments. After sorting, the initial vector or subspace for the existing algorithms was set to the eigenvectors from the previous problem (the modified version ’*’). We compared the computational time across different methods. All experiments were conducted on the Helmholtz operator dataset, with a matrix dimension of 6400 and a tolerance of 1e-8. The results are shown in Table 2.
@@ 4.4 Efficacy of Sorting Algorithms
We analyze the performance of the sorting algorithm module from two perspectives: 1. Comparing the performance of SCSF algorithm with and without ’sorting’ as shown in Table 3. 2. Evaluating the effectiveness of different sorting algorithms as detailed in Tables 4 and 5. We note that if the setting is ’w/o sort’, SCSF is approximately equivalent to directly using the Chebyshev subspace filter. Unlike the ChFSI used in the main experiments, the initialization of each solve in the ’w/o sort’ SCSF is set based on the information obtained from solving the previous problem (following the default unsorted sequence).
@@
Firstly, Table 3 indicates that incorporating sorting can improve SCSF speed to 1.3 to 2.8 times, reduce the number of iterations by 5% to 50%, and decrease total Flops by 7% to 43%. The effect of sorting is more pronounced with smaller numbers of solutions $L$. This is because when $L$ is large, the inherited subspace already contains most of the necessary correlation information, diminishing the impact of sorting. Moreover, the Flops in the Filter component constitute over 70% of SCSF’s computational load. A detailed time analysis of different aspects of SCSF can be found in Appendix E.3. Additionally, the ’w/o sort’ SCSF achieves a computational speedup of 1.2 to 1.5 times compared to the ChFSI used in the main experiments. The primary difference lies in their initialization strategies: ChFSI uses random initialization for each solve, whereas the ’w/o sort’ SCSF leverages information from the previous problem for initialization. This indicates that, even without sorting, there is a certain level of similarity between problems in the dataset. Such similarity can effectively accelerate the solving process.
@@
Secondly, as shown in Table 4, our designed truncated FFT sorting algorithm incurs significantly lower time cost compared to the complete greedy sorting in SKR (Wang et al., 2024), with its benefits becoming more pronounced as the dataset size increases. In the truncated FFT sorting algorithm, the FFT contributes minimally to computational overhead but significantly reduces the time required for subsequent greedy sorting. Table 5 shows SCSF solution times for matrices sorted using either greedy or truncated FFT sorting are nearly identical, highlighting its effectiveness.
Furthermore, our experiments show that as long as the truncated FFT is configured with reasonable parameters (e.g., truncating at $p_{0}=20$, where $p_{0}$ is much smaller than the dimension $p$ of the parameter matrix $P$), it achieves excellent performance without the need for a large $p_{0}$. For related experiments, please refer to Appendix E.4.3.
@@ 5 Conclusions and Future Work
In this paper, we addressed the critical bottleneck of generating large-scale eigenvalue datasets for training neural operators. We introduced SCSF, the first method to accelerate eigenvalue dataset generation by exploiting operator similarity. By integrating a truncated FFT sorting algorithm with a Chebyshev subspace filter, SCSF transforms the generation task into an efficient sequential solving problem. Our method achieves up to a 3.5x speedup over traditional solvers, significantly reducing computational redundancy. By lowering a key barrier to entry, SCSF provides a valuable tool for advancing research in the AI for Science community. For future work, we identify two primary directions. First, extending SCSF to handle nonlinear eigenvalue problems would broaden its applicability. Second, developing more effective distance metrics for the sorting algorithm could further improve its efficiency by better capturing operator similarity.
@@ B.1 Eigenvalue Datasets and Neural Eigenvalue Methods
Eigenvalue datasets are widely utilized in neural eigenvalue methods. In molecular chemistry research, eigenvalue algorithms are commonly employed to determine critical molecular properties, such as orbital energy levels (Kittel & McEuen, 2018). These properties form the foundation of datasets and are obtained by solving the eigenvalue problem of the Schrödinger equation and the Hamiltonian operator (Helgaker et al., 2013). Prominent datasets in this domain include QM7 (Blum & Reymond, 2009), QM9 (Ramakrishnan et al., 2014), ANI-1 (Smith et al., 2017), and MD17 (Chmiela et al., 2017). In materials science, eigenvalue algorithms are often applied to solve for electronic band structures and density of states in materials. Representative datasets in this field include the materials project (Jain et al., 2020) and OQMD (Kirklin et al., 2015). These datasets have been extensively used to train and validate neural eigenvalue methods (Schütt et al., 2017; Bartók et al., 2017; Rupp et al., 2012), driving advancements in molecular property prediction and materials design. In fluid dynamics and structural mechanics, eigenvalue algorithms are frequently utilized for modal analysis. Recently, many data-driven modal analysis algorithms have emerged, requiring eigenvalue datasets corresponding to differential operators for training (Murata et al., 2020; Iwata & Kawahara, 2023; Alford-Lago et al., 2022; Brunton et al., 2020; Akkad & He, 2023). Additionally, some studies leverage operator eigenvalue datasets to optimize algorithms. For instance, Luo et al. accelerates the solution of linear systems by predicting the eigenfunctions of operators.
@@ B.2 Eigenvalue Data Generation Algorithms
Nonetheless, traditional methods were not designed for dataset generation, resulting in high computational costs, which have become a significant barrier to the advancement of data-driven approaches (Zhang et al., 2023; Hao et al., 2022). Recent data augmentation research (Brandstetter et al., 2022; Liu et al., 2023) has led to the development of methods that preserve symmetries and conservation laws, enhancing model generalization and data efficiency. Wang et al. (2024); Dong et al. (2024) report acceleration in the process of solving linear equations, thereby speeding up the generation of PDE datasets.

### v2

@@ 1 Introduction
Despite their success, data-driven approaches face a fundamental limitation: the reliance on labeled datasets. Training neural networks requires large-scale labeled data, which is often generated using computationally expensive traditional methods. It usually takes dozens of hours or even days. For example, the QM9 dataset (Ramakrishnan et al., 2014) contains $1.34\times 10^{5}$ molecular data points, each produced by solving Hamiltonian operator eigenvalue problems. These calculations typically employ traditional algorithms, whose computational costs can escalate dramatically with increasing problem complexity, like finer grid resolutions or higher accuracy requirements. This scalability issue represents a significant bottleneck for generating the labeled data needed to train deep learning models. Furthermore, the diversity of scientific problems leads to the need for a unique dataset for each scenario, which further intensifies this challenge of computational intractability. As a result, the high computational expense of generating eigenvalue data severely limits the application of data-driven approaches (Ren et al., 2026; Zhang et al., 2023; Huang et al., 2025; Liu et al., ; Liu et al., 2024; 2025c; Wu et al., 2025; Luo et al., 2024; Liu et al., 2025b; Zhang et al., 2024; Wang et al., a; Hou et al., 2026; Zhang et al., 2026).
@@
  • To the best of our knowledge, SCSF is the first method to accelerate operator eigenvalue dataset generation by exploiting intrinsic operator similarity through a novel sorting and filtering framework.
@@ 2.1 Discretization of Eigenvalue Problem
This section outlines the discretization of operator eigenvalue problems into matrix form. Our method focuses on solving these matrix problems, which is the most time-consuming step in generating the required eigenvalue datasets. As shown in Figure 1 (left), these problems are typically solved by numerical discretization methods such as FDM (Strikwerda, 2004; LeVeque, 2002). These discretization techniques embed the infinite-dimensional Hilbert space of operators into an appropriate finite-dimensional space, thereby transforming operator eigenvalue problems into matrix eigenvalue problems. We provide a simple example to clarify the discussed processes. A detailed process can be found in Appendix C. Specifically, we discuss the case that uses FDM to solve the eigenvalue problem of the two-dimensional Poisson operator, transforming it into a matrix eigenvalue problem:
@@
Typically, training a neural network requires a number of data from $10^{3}$ to $10^{5}$ (Lu et al., 2019). Such a multitude of eigenvalue systems, derived from the same distribution of operators, naturally exhibit a high similarity (Soodhalter et al., 2020). It is precisely this similarity that is key to the effective acceleration of SCSF. We can conceptualize this as the task of solving a sequential series of matrix eigenvalue problems:
$A^{(i)}v^{(i)}_{j}=\lambda^{(i)}_{j}v^{(i)}_{j},\quad j=1,\cdots,L;\quad i=1,2,\cdots,N$ (3)
@@ 4 $\sigma_{i+1}=1/(2/\sigma_{1}-\sigma_{i})$;
This section details the Chebyshev filter, which is constructed using Chebyshev polynomials. Distinguished by their optimal uniform approximation properties and efficient three-term recurrence relations, these are among the most widely utilized orthogonal polynomials (Mason & Handscomb, 2002; Rivlin, 2020). This filtering technique is particularly effective for our method because it can efficiently construct a polynomial that isolates a target spectral interval, allowing us to use the solution of one eigenvalue problem to significantly accelerate the next.
The Chebyshev polynomials $C_{m}(t)$ of degree $m$ are defined on [1,1][-1,1] and are expressed as
@@ 3.1 The Sorting Algorithm
To benefit the successive solving sequence of the eigenvalue problem, we need a sorting algorithm that pulls matrices with similar spectral properties, like invariant subspaces, close enough in the solving sequence, so that solving the current matrix in sequence can be easily boosted by the previous solving. Recalling Section 2.1, the matrix $A^{(i)}$ is the operator matrix derived through numerical discretization (e.g., Finite Difference Method) from the parameter matrix $P^{(i)}$ (e.g., the coefficient $k(x,y)$ in the Poisson equation) (Lu et al., 2022; Li et al., 2020). For instance, in the Poisson equation, the values $k_{i,j}$ from $P^{(i)}$ become part of the discretized sparse matrix $A^{(i)}$, as shown in Eq. 2. A naive strategy is to use the Frobenius distance of the parameter matrices $P^{(i)}$ to perform a greedy sort (Wang et al., 2024). And by repeatedly fetching without reservation from the remaining matrix in the dataset, we can reorganize the solving sequence so that the successive solving can benefit from the re-ordered sequence.
@@ 12  Remove $j_{min}$ from $seq_{0}$, append $j_{min}$ to ${seq}_{mat}$ and set $i_{0}=j_{min}$;
As shown in Algorithm 2, suppose we have $N$ eigenvalue problems, the parameter matrices $P^{(i)}\in\mathds{C}^{p\times p}$, and the low-frequency truncated matrices $P^{(i)}_{low}\in\mathds{C}^{p_{0}\times p_{0}}$. The computational complexity of directly using a greedy algorithm is $\mathcal{O}(N^{2}p^{2})$. Our sorting algorithm’s complexity consists of two main parts: 1. FFT Computation: The complexity of FFT is $\mathcal{O}(p^{2}\log p)$ per matrix. For $N$ matrices, this totals $\mathcal{O}(Np^{2}\log p)$. 2. Greedy Sorting: The subsequent greedy sorting algorithm has a complexity of $\mathcal{O}(N^{2}p_{0}^{2})$. Overall, the total complexity is $\mathcal{O}(N^{2}p_{0}^{2}+Np^{2}\log p)$. Since $p_{0}\ll p$ and $p\ll N$, our sorting algorithm effectively reduces computational cost. In addition, we conducted a qualitative theoretical analysis of the truncated FFT to demonstrate its rationality. Details can be found in Appendix F.
@@ 3.2 Chebyshev Filtered Subspace Iteration
Algorithm 3 outlines the process of ChFSI for solving the $i$-th eigenvalue problem $A^{(i)}$ ($1<i\leqslant N$) where $L$ eigenvalues need to be solved. The initial approximate invariant subspace $V^{(i-1)}$ and spectral distribution $\Lambda^{(i-1)}$ are derived from the eigenvectors and eigenvalues of the previous problem $A^{(i-1)}$ in the sequence. This initialization strategy, illustrated in Figure 2 (g), acts as a “warm start”, where a high-quality initial subspace significantly reduces the computational cost of subsequent iterations. The parameter $m$ denotes the polynomial degree in the filter function, e.g., $m=20$. As depicted in Figure 2 (f), the Chebyshev filter modifies the operator’s spectrum to amplify the target region, effectively transforming the problem into an equivalent one that is easier to solve. For the first eigenvalue problem $A^{(1)}$ in the sequence, the initial iterative subspace $\tilde{V}_{0}$ and initial spectrum $\tilde{\Lambda}_{0}$ are randomly generated.
@@ 4.2 Main Experiment
Table 1: Comparison of average computation times (in seconds) for eigenvalue problems using various algorithms. The first row lists different algorithms, the first column details the datasets, including matrix dimensions and solution precisions (relative residual), and the second column shows the number of eigenvalues $L$ computed for each matrix. The best algorithm is in bold. The symbol ‘-’ denotes the result of a method that fails to converge under the given setting.
@@ 4.3 Efficacy of Chebyshev Subspace Filter
Table 2: Impact of initial subspace modifications on average computation time (in seconds) for different algorithms. ‘*’ denotes the modified version. The first row lists algorithms, and the first column shows the number of eigenvalues $L$ computed. The best algorithm is in bold, and ‘-’ indicates the result of a method that fails to converge under the given setting.
@@
To analyze the efficacy of the Chebyshev subspace filter, we conducted the following experiments. After sorting, the initial vector or subspace for the existing algorithms was set to the eigenvectors from the previous problem (the modified version ‘*’). We compared the computational time across different methods. All experiments were conducted on the Helmholtz operator dataset, with a matrix dimension of 6400 and a tolerance of 1e-8. The results are shown in Table 2.
@@ 4.4 Efficacy of Sorting Algorithms
We analyze the performance of the sorting algorithm module from two perspectives: 1. Comparing the performance of SCSF algorithm with and without ‘sorting’ as shown in Table 3. 2. Evaluating the effectiveness of different sorting algorithms as detailed in Tables 4 and 5. We note that if the setting is ‘w/o sort’, SCSF is approximately equivalent to directly using the Chebyshev subspace filter. Unlike the ChFSI used in the main experiments, the initialization of each solve in the ‘w/o sort’ SCSF is set based on the information obtained from solving the previous problem (following the default unsorted sequence).
@@
Firstly, Table 3 indicates that incorporating sorting can improve SCSF speed to 1.3 to 2.8 times, reduce the number of iterations by 5% to 50%, and decrease total Flops by 7% to 43%. The effect of sorting is more pronounced with smaller numbers of solutions $L$. This is because when $L$ is large, the inherited subspace already contains most of the necessary correlation information, diminishing the impact of sorting. Moreover, the Flops in the Filter component constitute over 70% of SCSF’s computational load. A detailed time analysis of different aspects of SCSF can be found in Appendix E.3. Additionally, the ‘w/o sort’ SCSF achieves a computational speedup of 1.2 to 1.5 times compared to the ChFSI used in the main experiments. The primary difference lies in their initialization strategies: ChFSI uses random initialization for each solve, whereas the ‘w/o sort’ SCSF leverages information from the previous problem for initialization. This indicates that, even without sorting, there is a certain level of similarity between problems in the dataset. Such similarity can effectively accelerate the solving process.
Secondly, as shown in Table 4, our designed truncated FFT sorting algorithm incurs significantly lower time cost compared to the complete greedy sorting in SKR (Wang et al., 2024), with its benefits becoming more pronounced as the dataset size increases. In the truncated FFT sorting algorithm, the FFT contributes minimally to computational overhead but significantly reduces the time required for subsequent greedy sorting.
@@
Table 5 validates the accuracy of our efficient sorting against the expensive greedy approach. To ensure fairness, we averaged the results over 100 independent runs to eliminate fluctuations. The results for ‘Greedy’ and ‘Ours’ are identical because the sequences produced by both algorithms are over 98% identical. Consequently, the downstream solver incurs exactly the same operation cost, demonstrating that our low-cost sorting achieves the same acceleration performance as the high-cost greedy approach.
Furthermore, our experiments demonstrate that the choice of the truncation threshold, $p_{0}$, is robust. We found that excellent performance is achieved with a relatively small $p_{0}$ (e.g., $p_{0}=20$), a value much smaller than the parameter matrix dimension $p$. This is because performance stabilizes once a sufficient number of low-frequency components are captured, making it unnecessary to retain a large number of components. For related experiments, please refer to Appendix E.4.3.
@@ 5 Conclusions and Future Work
Conclusions
[…]
Limitations and Future Work
Despite SCSF’s demonstrated efficiency, we have identified three key directions for future development. First, we aim to extend the current framework from Hermitian linear operators to non-Hermitian and non-linear operators. Second, the correlation between operators is currently measured using the Frobenius norm of their parameter matrices. We plan to explore more effective, problem-specific metrics rather than relying on a fixed one. Finally, SCSF’s effectiveness hinges on a continuous mapping between parameters and operators, rendering it ineffective when this mapping is discontinuous. Future work will involve measuring correlation directly from the operators or their matrix representations to bypass the challenges of discontinuous parameterization.
Acknowledgements
The authors would like to thank all the anonymous reviewers for their insightful comments and valuable suggestions. This work was supported by the National Key R&D Program of China (No. 2022ZD0119801), the National Natural Science Foundation of China (Nos. U23A20388, 62021001), and the Smart-Grid National Science and Technology Major Project (No. 2025ZD0805500).
@@ B.1 Eigenvalue Datasets and Neural Eigenvalue Methods
Eigenvalue datasets are widely utilized in neural eigenvalue methods. In molecular chemistry research, eigenvalue algorithms are commonly employed to determine critical molecular properties, such as orbital energy levels (Kittel & McEuen, 2018). These properties form the foundation of datasets and are obtained by solving the eigenvalue problem of the Schrödinger equation and the Hamiltonian operator (Helgaker et al., 2013). Prominent datasets in this domain include QM7 (Blum & Reymond, 2009), QM9 (Ramakrishnan et al., 2014), ANI-1 (Smith et al., 2017), and MD17 (Chmiela et al., 2017). In materials science, eigenvalue algorithms are often applied to solve for electronic band structures and density of states in materials. Representative datasets in this field include the materials project (Jain et al., 2020) and OQMD (Kirklin et al., 2015). These datasets have been extensively used to train and validate neural eigenvalue methods (Schütt et al., 2017; Bartók et al., 2017; Rupp et al., 2012), driving advancements in molecular property prediction and materials design. In fluid dynamics and structural mechanics, eigenvalue algorithms are frequently utilized for modal analysis. Recently, many data-driven modal analysis algorithms have emerged, requiring eigenvalue datasets corresponding to differential operators for training (Murata et al., 2020; Iwata & Kawahara, 2023; Alford-Lago et al., 2022; Brunton et al., 2020; Akkad & He, 2023; Wang et al., b). Additionally, some studies leverage operator eigenvalue datasets to optimize algorithms. For instance, Luo et al. accelerates the solution of linear systems by predicting the eigenfunctions of operators.
@@ B.2 Eigenvalue Data Generation Algorithms
Nonetheless, traditional methods were not designed for dataset generation, resulting in high computational costs, which have become a significant barrier to the advancement of data-driven approaches (Zhang et al., 2023; Hao et al., 2022; Chen et al., 2025). Recent data augmentation research (Brandstetter et al., 2022; Liu et al., 2023) has led to the development of methods that preserve symmetries and conservation laws, enhancing model generalization and data efficiency. Wang et al. (2024); Dong et al. (2024) report acceleration in the process of solving linear equations, thereby speeding up the generation of PDE datasets.